OpenStudy (anonymous):

if a spherical tank 2 m in diameter can be filled with a liquid for $300, find the cost to fill the tanks of 8 diameter.

OpenStudy (anonymous):

the area of the sphere tank with diameter 2m= 4.19m^2 (4/3pi.r^2) $300=4.19m^2 the are for sphere tank measuring 8m in diameter= 67.05m^2 67.05m^2/4.19m^2=16 16*300=$4800